OpenStudy (anonymous):

T*(RBJBR)=VPLNT Where the letters represent a digit in a number. "RBJBR" is five digits long. Find possible letter values. HOORAY FOR CRYPTO.

OpenStudy (anonymous):

@FoolForMath @KingGeorge @Mr.Math :P Get in here, whether you like it or not. @TuringTest

OpenStudy (anonymous):

any answer choices?

OpenStudy (anonymous):

No, lol, multiple choice is for dorks.

OpenStudy (anonymous):

lol what about fools?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Insane question!!

OpenStudy (anonymous):

I can solve it using a computer program, just traversing through all 5 digits palindromes.

OpenStudy (anonymous):

Dang it, @FoolForMath , you are such a jerk. If I wanted I could get Mathematica to solve this for me.

OpenStudy (anonymous):

6*(15251) = 91506 6*(15351) = 92106 6*(15451) = 92706 6*(15551) = 93306 6*(15651) = 93906 6*(15751) = 94506 6*(15851) = 95106 6*(15951) = 95706 6*(16061) = 96366 6*(16161) = 96966 6*(16261) = 97566 6*(16361) = 98166 6*(16461) = 98766 6*(16561) = 99366 6*(16661) = 99966 7*(10001) = 70007 7*(10101) = 70707 7*(10201) = 71407 7*(10301) = 72107 7*(10401) = 72807 7*(10501) = 73507 7*(10601) = 74207 7*(10701) = 74907 To name a few. Wrote a program.

OpenStudy (anonymous):

Please be respectful. There are many solution to this problem so, the problem in this form is just not worth for a manual attempt.

OpenStudy (anonymous):

>:| Maybe I should tell that to the math dept. then?

OpenStudy (anonymous):

We are expect to find possible solutions by hand.

OpenStudy (anonymous):

Not all of them, mind you.

OpenStudy (anonymous):

0*(00000) = 00000

OpenStudy (anonymous):

They have to be unique, unfortunately.

OpenStudy (anonymous):

Forgot to mention. :(

OpenStudy (anonymous):

The digits? No problem.

OpenStudy (anonymous):

2*(15451) = 30902 2*(18481) = 36962 3*(15951) = 47853 3*(16261) = 48783 3*(16761) = 50283 3*(16961) = 50883 3*(18081) = 54243 3*(19091) = 57273 3*(19291) = 57873 4*(12721) = 50884 4*(15751) = 63004 4*(15951) = 63804 4*(17071) = 68284 4*(17271) = 69084 4*(18081) = 72324 4*(19091) = 76364 4*(19591) = 78364 5*(14641) = 73205 5*(16461) = 82305 5*(17271) = 86355 5*(18681) = 93405 6*(13431) = 80586 6*(14541) = 87246 6*(15451) = 92706 7*(12421) = 86947 7*(13631) = 95417 7*(14041) = 98287

OpenStudy (anonymous):

6*(15451) = 92706 is a possible solution. There could be more!

OpenStudy (anonymous):

Ah, I'm not looking for guessed numbers, or algorithmically determined ones. I'm looking for a formally written proof that logically deduces possible values.

OpenStudy (kinggeorge):

Well, it's easy to find R. Since \(T\cdot R =T\mod 10\) it must be true that \(R=1\).

OpenStudy (anonymous):

T * R = T mod 10 2 * 6 = 2 mod 10

OpenStudy (kinggeorge):

Good point. T=5, R=3 also works.

OpenStudy (anonymous):

Weird how 1 is the only of value of R that works though, huh.

OpenStudy (anonymous):

(That is, unless I messed up writing my program, but the list with unique digits that I posted was the entire output)

OpenStudy (kinggeorge):

It is still true that R=1 however. You just need to look at the other side. Since RBJBR is a palindrome, if R is greater than 1 such that \(TR=T\mod 10\), then \(TR\geq 10\) Thus, if \(R\geq1\) The number VPLNT would have 6 digits. We conclude that \(R=1\)

OpenStudy (anonymous):

Ahh, excellent point.

OpenStudy (kinggeorge):

If I'm doing this correct, we also have that \(T\cdot B\equiv N \mod 10\)

OpenStudy (kinggeorge):

Since these all values are distinct, this means that T and B are non-zero.

OpenStudy (kinggeorge):

More strongly, \(B, T \geq 2\)

OpenStudy (kinggeorge):

Another fact we know: \(L= T\cdot J\mod 10\) plus the second digit of \(T\cdot B\) taken mod 10.

OpenStudy (kinggeorge):

Also, \(V>T\) so \(T\leq 8\).

OpenStudy (kinggeorge):

Let's list some pairs \((T, B)\) that are impossible with our current criteria. (Ignoring those less than 2 or those where T=B) (2, 6) (6, 2) (4, 6) (6, 4) (6, 8) (8, 6) (5, 3) (3, 5) (5, 7) (7, 5) (5, 9) (9, 5) (3, 7) (7, 3)

OpenStudy (kinggeorge):

That leaves us with the possible pairs of (we can also reverse these) (2, 3) (2, 4) (2, 5) (2, 7) (2, 8) (2, 9) (3, 4) (3, 6) (3, 8) (3, 9) (4, 5) (4, 7) (4, 8) (4, 9) (5, 6) (5, 8) (6, 7) (6, 9) (7, 8) (7, 9) (8, 9) Let's see if we can narrow this down a bit. :/

OpenStudy (anonymous):

I feel a bit like a jerk not participating in the problem solving process, but when you're on, I'm not, and vice versa. And now I need to go to sleep. :(

OpenStudy (kinggeorge):

Good night. I'm sort of at a loss of where to go from here.

OpenStudy (anonymous):

I'm perpetually at a loss.

OpenStudy (kinggeorge):

I suppose we could also say that P equals the sum of N and the second digit of \(T\cdot J\). I didn't show all steps to get that, but it's true.